OpenStudy (anonymous):

a store is offering a 25% discount on all items. Also employees get a 10% discount. if you are an employee, which discount would you want to be applied first to save the most money? 10% 25% it doesnt matter which discount is applied first the result is the same not enough info is given

OpenStudy (compassionate):

\(\Huge\bf Welcome~to~ \color{#00B4ff}{Open}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge\bf\color{#7cc517}{Study~!!}~\) both lead to the same final price here's why let x = original price Lets say you do the 25% discount first then the 10% discount apply the first discount x - 0.25x = 0.75x Then apply the second discount 0.75x - 0.10*0.75x = 0.675*x So in the end, you're paying 67.5% of the original price and you get an overall discount of 100 - 67.5 = 32.5% off ---------------------------------------------------------------- Now let's say you do the 10% discount first then the 25% discount apply the first discount x - 0.10x = 0.90x then apply the second discount 0.90x - 0.25*0.90x = 0.675*x and you get the same final price So either way, you're getting the same discount (no matter which one you do first) The two discounts basically combine to 32.5% discount in the end. - Courtesy of Jim Johnson. Thank you for visiting! Come back! I'll be glad to help you! !! :)

OpenStudy (anonymous):

Thank you that helped alot i understand now!